OpenStudy (anonymous):

f(x)= 8 - 0.5x^2 on the interval [-1,4] What is the actual area under the curve using a right sum with 10 rectangles? (Solve using limits) I'm not sure how to set this up.. any ideas?

OpenStudy (jennychan12):

\[\Delta x = \frac{ 5 }{ 10 } = \frac{ 1 }{ 2 }\] \[x _{i} = -1+\frac{ i }{ 2 }\] \[\frac{ 1 }{ 2 }\sum_{i=1}^{10} [8-0.5(-1+0.5 (\frac{ i }{ 2 })^2]\] \[\frac{ 1 }{ 2 } \sum_{i=1}^{10} [8-0.5(-1+\frac{ 1 }{ 4 } (i^2)\] \[\frac{ 1 }{ 2 } \sum_{i=1}^{10} [8-0.5(-1+\frac{ 1 }{ 4 } (\frac{ n(n+1)(2n+1) }{ 6 })\] plug in 1-10 for i add together